Harry deposits $2000 in a saving account that compounds annually, and plans to leave the money in the account for 2 years. If Harry wants his account to increase in value to $3300, then the interest rate must be greater than what percent?

Answer by lwsshak3(11628)

Harry deposits $2000 in a saving account that compounds annually, and plans to leave the money in the account for 2 years. If Harry wants his account to increase in value to $3300, then the interest rate must be greater than what percent?
***
Compound Interest Formula: A=P(1+i)^n, P=initial investment, i=interest rate per compounding period, n=number of periods, A=amount after n-periods.
For given problem:
P=2000
i=to be determined
n=2
A=3300
..
3300=2000(1+i)^2
3300/2000=(1+i)^2
33/20=(1+i)^2
take sqrt of both sides
√(33/20)=1+i
i=√(33/20)-1=0.2845 (rounded)
interest rate must be greater than what percent? 28.45%